OpenStudy (anonymous):

What is the solution to the equation 36^3c = square root of 6^(8c – 6) ?

hero (hero):

Did you square both sides yet?

OpenStudy (anonymous):

square root of 36 is 6 and the square root of 6 is 2.4494897?

hero (hero):

It doesn't exactly work like that You can't take the square root of the 6 while it has an exponent attached to it

hero (hero):

You have to square both sides to get rid of the square root.

OpenStudy (anonymous):

can you walk me through this please?

hero (hero):

I'm trying, but LaTeX is not working too well right now

hero (hero):

\[36^{3c} = \sqrt{6^{(8c – 6)}}\]

hero (hero):

That's what you probably started off with

hero (hero):

Now you have to square both sides like this: \[{(36^{3c})}^2 = ({\sqrt{6^{(8c – 6)}}})^2\]

OpenStudy (anonymous):

yes

hero (hero):

The square and the square root cancels to get this: \[{(36^{3c})}^2 =6^{(8c – 6)}\]

hero (hero):

Meanwhile on the left side of the equation, you multiply exponents 3c and 2 to get: \[36^{(6c)} =6^{(8c – 6)}\]

hero (hero):

Now, 36 can be expressed as \(6^2\), therefore we re-write the equation producing: \[\large6^{2^{(6c)}} =6^{(8c – 6)}\]

hero (hero):

But since both 2 and 6c are exponents we can express it as a multiplication producing: \[6^{2{(6c)}} =6^{(8c – 6)}\]

hero (hero):

Now since the bases on both sides of the equation are the same we can simply set the exponents on both sides equal to one another producing: 2(6c) = 8c - 6 Can you solve for c from here?

OpenStudy (anonymous):

no Im still confused

hero (hero):

Maybe reviewing rules of exponents will help

OpenStudy (anonymous):

okay I tried working itout and I got 4/3?

hero (hero):

Weird. I got something else.

OpenStudy (anonymous):

can you work it out please?

hero (hero):

2(6c) = 8c - 6 12c = 8c - 6 12c - 8c = -6 c(12-8) = -6 c(4) = -6 c = -6/4 c = -3(2)/2(2) c = -3/2

OpenStudy (anonymous):

okay, I see what I did wrong. Thank you